# CosmoBench: A Multiscale, Multiview, Multitask Cosmology Benchmark for Geometric Deep Learning

### v1

@@ B.2 Predicting Velocities from Positions
Following the discussion Sec. 4.2, the linearized continuity equation implies that
@@
where $a$ is the scale factor, $H=\dot{a}/a$ is the Hubble parameter, $f$ is the linear growth rate (which is a function of the cosmological parameters). Furthermore, in linear theory, the overdensity of halos (or galaxies) $\delta_{h}$ is related to the overdensity of matter $\delta$ via linear rescaling by a constant $b$, such that $\delta_{h}=b\,\delta$. Equation 11 can thus be written in Fourier space, yielding the velocity in Fourier space $\bm{v}(\bm{k})$, as
@@
While Equation Equation 3 provides an integral solution in position space (i.e. $\bm{x}$ space), the Fourier exposition of Equation Equation 12 provides a practical, commonly used, way to infer the velocity field from the density field. We first compute $\delta_{h}(\bm{x})$ by computing the halo density on a grid of size $N_{g}$ (effectively by counting the number of halos in each cell, and interpolating using a triangular-shaped cloud scheme). We then Fourier transform to obtain $\delta_{h}(\bm{k})$. The value of $\bm{k}$ on each Fourier grid point is given as $2\pi/L(i,j,k)$, where $i,j,k\in\{0,1,...N_{g}-1\}$. We choose the value of $N_{g}$ that gives the best $R^{2}$ on the velocity prediction for the validation set, considering values between $N_{g}=10$ and $90$ in increments of $10$. Up to the proportionality constant, we evaluate $\bm{v}(\bm{k})$ and then Fourier transform back to obtain $\bm{v}(\bm{x})$. All these computations were performed using the Pylians package [^4] https://pylians3.readthedocs.io/.
@@
To visualize the predicted velocities of the point clouds, as in Fig. 2, we compute the velocities on a 3D grid, and then project along one dimension. We use the yt package [^5] https://yt-project.org/. Fig. 2 shows results for Quijote, with the true velocities in the left panel, followed by the linear-theory oracle, LLS, and GNN from left to right. We observe that the true velocity field contains many complex flow paths, while the linear theory and LLS predictions tend to predict smoother flows, typically following the large-scale bulk motion. This is expected due to their linear nature. The GNN also produces smoother predictions than the truth, but it picks up on some of the more smaller-scale flows. Take Fig. 2 as an example: the GNN predicts a small local sink towards the lower middle of the panel, while the linear predictions have all flow lines going in the same direction towards a larger sink.
@@
For further visualization, Fig. B.3 provides a similar plot to Fig. 2, but for two different test set point-cloud examples from each of the multiscale simulations (Quijote, CAMELS-SAM, and CAMELS). We additionally report the $R^{2}$ for the specific point cloud being plotted in the top right corner of each panel. The trends for the two Quijote examples (first and fourth rows) are similar to Fig. 2, although it is interesting to note that the fourth row is an example where LLS is better than the GNN. CAMELS-SAM and CAMELS produce much smoother true velocities in the image, as they are far smaller scale systems than Quijote, thus the chosen grid resolution traces the velocities more smoothly. For CAMELS-SAM, the two examples (second and fifth rows) are consistent with the average behavior reported in Tab. 3, with the GNN performing best, while for CAMELS the linear theory oracle performs the best. The most accurate performance of the linear theory oracle for CAMELS is a notable result, as linear theory is known to break down on the small scales probed by CAMELS, but nonetheless, this implies that a linear relation is sufficient to reasonably predict the velocity (more accurately so compared to the other methods currently in this benchmark) if one exactly knew the constant of proportionality, as our oracle does. A future entry to the benchmark could be to use linear, or even non-linear, theory without any oracle-like information by performing the non-trivial task of jointly inferring the density, cosmological parameters, and in turn the velocity via Bayesian hierarchical modeling (see e.g. [66, 67] for works in this direction).
@@ C.1 Predicting Cosmological Parameters from Merger Trees
In this section we provide further details about the 1-Nearest Neighbor predictor introduced in Sec. 5.1. We consider for each tree the empirical distribution across nodes of three features: the concentration $c$, the maximum velocity $v_{\text{max}}$, and the scale factor $a$ (a measure of time). Consider two empirical distributions, $\widehat{P}_{A}$ and $\widehat{P}_{B}$ of a univariate random variable $X$. The Kolmogorov-Smirnov statistic [e.g. 69] is
@@
As shown in Tab. 4 (left), $c$ and $v_{\rm max}$ are the most informative features when inferring $\Omega_{m}$, while $a$ is most informative for $\sigma_{8}$. Combining all nodes features achieves far superior accuracy. To examine the model performance further, we present scatter plots of the target cosmological parameters ($x$-axis) and the predicted parameters ($y$-axis) from (a) DeepSet and (b) GNN: Fig. C.1 shows the results from using the node feature scale factor $a$ only, whereas Fig. C.2 shows the results using all features (mass $M$, concentration $c$, halo maximum circular velocity $v_{\max}$, and scale factor $a$). We see mild improvements of GNN over DeepSet; both models are highly predictive of $\Omega_{\rm m}$ when using all features. In this work, we make use of the entire merger tree from CS-Trees to predict cosmological parameters; using a subtree for such prediction task is an interesting direction for future research (e.g. the main branch used for generative modelling tasks in [85]).
@@ C.2 Reconstructing Fine-Scale Merger Trees
We adopt and extend the EPS implementation from [73] [^6] https://github.com/shergreen/SatGen. The original algorithm is intended to generate many merger trees in parallel for a given set of constraints $s=\{\Omega_{\rm m},\sigma_{8},\ldots\}$, which include cosmological parameters and halo conditions (e.g., the range of root halo mass, the minimum mass threshold of the leaf halos). For our application in classifying unresolved merger nodes, each virtual node $v\in\tilde{V}$ in the coarsened tree $\mathcal{T}_{c}$ (see Algorithm 1) defines a specific constraint set $s_{v}$, derived from its post-merger node and pre-merger nodes. We collect a list of constraints $\mathbf{s}=[s_{v(1)},s_{v(2)},\ldots]$ induced from $20\%$ of virtual nodes across $n_{\mathcal{T}}=97$ coarsened trees, and then iterate through each constraint $s\in\mathbf{s}$ to generate $5$ merger trees using the EPS algorithm. The sequential processing over all constraints, of total size $|\mathbf{s}|=O(n_{\mathcal{T}}|\tilde{V}|)$, is currently the main computational bottleneck. We note that this step could be significantly accelerated by parallelizing over constraints, which we leave for future work.
@@
As reported in Tab. 4 (right), the ML approaches perform equivalently or better than the cosmology baseline (EPS). Moreover, they require considerably less time. Nevertheless, all baselines considered still have room for improvement. As a first step, we only consider predicting the unresolved merger node labels (i.e. whether there exists a merger node or not). Another important future step is to also predict the unresolved merger node features, which would enable enhanced inference quality from incomplete or coarse-grained merger trees.

### v2

@@ B.2 Predicting Velocities from Positions
Following the discussion Sec.˜4.2, the linearized continuity equation implies that
@@
where $a$ is the scale factor, $H=\dot{a}/a$ is the Hubble parameter, $f$ is the linear growth rate (which is a function of the cosmological parameters). Furthermore, in linear theory, the overdensity of halos (or galaxies) $\delta_{h}$ is related to the overdensity of matter $\delta$ via linear rescaling by a constant $b$, such that $\delta_{h}=b\,\delta$. Equation˜11 can thus be written in Fourier space, yielding the velocity in Fourier space $\bm{v}(\bm{k})$, as
@@
While Equation Equation˜3 provides an integral solution in position space (i.e. $\bm{x}$ space), the Fourier exposition of Equation Equation˜12 provides a practical, commonly used, way to infer the velocity field from the density field. We first compute $\delta_{h}(\bm{x})$ by computing the halo density on a grid of size $N_{g}$ (effectively by counting the number of halos in each cell, and interpolating using a triangular-shaped cloud scheme). We then Fourier transform to obtain $\delta_{h}(\bm{k})$. The value of $\bm{k}$ on each Fourier grid point is given as $2\pi/L(i,j,k)$, where $i,j,k\in\{0,1,...N_{g}-1\}$. We choose the value of $N_{g}$ that gives the best $R^{2}$ on the velocity prediction for the validation set, considering values between $N_{g}=10$ and $90$ in increments of $10$. Up to the proportionality constant, we evaluate $\bm{v}(\bm{k})$ and then Fourier transform back to obtain $\bm{v}(\bm{x})$. All these computations were performed using the Pylians package [^5] https://pylians3.readthedocs.io/.
@@
LLS Ablation
We find the results of LLS depends crucially on using the ensembles of features (inverse of powers of pairwise distances for all power less or equal to $P$). In Tab.˜10 we run LLS with a single choice of power order $p$ at a time (where $K=10$). We see that the performance of LLS on a single power order is not superior to GNN.
Table 10: The Effect of LLS Features for Velocity Prediction on Quijote
[…]
$\mathbf{R^{2}}$ $\uparrow$ 0.408 0.404 0.385 0.000 0.435 0.410
@@
To visualize the predicted velocities of the point clouds, as in Fig.˜2, we compute the velocities on a 3D grid, and then project along one dimension. We use the yt package [^6] https://yt-project.org/. Fig.˜2 shows results for Quijote, with the true velocities in the left panel, followed by the linear-theory oracle, LLS, and GNN from left to right. We observe that the true velocity field contains many complex flow paths, while the linear theory and LLS predictions tend to predict smoother flows, typically following the large-scale bulk motion. This is expected due to their linear nature. The GNN also produces smoother predictions than the truth, but it picks up on some of the more smaller-scale flows. Take Fig.˜2 as an example: the GNN predicts a small local sink towards the lower middle of the panel, while the linear predictions have all flow lines going in the same direction towards a larger sink.
@@
For further visualization, Fig.˜B.3 provides a similar plot to Fig.˜2, but for two different test set point-cloud examples from each of the multiscale simulations (Quijote, CAMELS-SAM, and CAMELS). We additionally report the $R^{2}$ for the specific point cloud being plotted in the top right corner of each panel. The trends for the two Quijote examples (first and fourth rows) are similar to Fig.˜2, although it is interesting to note that the fourth row is an example where LLS is better than the GNN. CAMELS-SAM and CAMELS produce much smoother true velocities in the image, as they are far smaller scale systems than Quijote, thus the chosen grid resolution traces the velocities more smoothly. For CAMELS-SAM, the two examples (second and fifth rows) are consistent with the average behavior reported in Tab.˜3, with the GNN performing best, while for CAMELS the linear theory oracle performs the best. The most accurate performance of the linear theory oracle for CAMELS is a notable result, as linear theory is known to break down on the small scales probed by CAMELS, but nonetheless, this implies that a linear relation is sufficient to reasonably predict the velocity (more accurately so compared to the other methods currently in this benchmark) if one exactly knew the constant of proportionality, as our oracle does. A future entry to the benchmark could use linear, or even non-linear, theory without any oracle-like information by performing the non-trivial task of jointly inferring the density, cosmological parameters, and in turn the velocity via Bayesian hierarchical modeling (see e.g. [66, 67, 83, 84] for works in this direction).
B.3 Predicting Velocities from Redshift Positions
Redshift Data Details
In cosmology, the redshift position vector $\bm{s}\in{\mathbb{R}}^{3}$ is obtained by displacing the real-space position $\bm{x}\in{\mathbb{R}}^{3}$ along the line of sight proportional by the object’s peculiar velocity,
$\bm{s}=\bm{x}+\frac{\bm{v}^{\top}\hat{\bm{n}}}{aH}\hat{\bm{n}}$, (13)
[…]
Figure B.4: Comparison between real-space positions (left panel) and redshift positions (right panel).
Linear Theory Modification Details
Recall in real space, linear theory relates the overdensity of halos (or galaxies) $\delta_{h}$ to the overdensity of matter $\delta$ linearly, $\delta_{h}=b\,\delta$. In redshift space this is modified by the Kaiser effect [85], such that $\delta_{h}=(b+f\mu^{2})\delta$, where $f\approx\Omega_{\rm m}^{0.55}$ and $\mu$ is the cosine of the angle to the axis of redshift distortion (here the $z$-axis). This modification essentially stretches the field to account for the redshift-space distortions. For our linear theory oracle we treat $b$ and $f$ as free parameters which are fit to the data.
LLS Modification Details
For the modified variant of LLS, we scale the $z$-axis distance by a shrinkage factor $\lambda$, where $\lambda$ is searched over $\{0.95,0.99\}$ for Quijote, and $\{0.3,0.4,0.5,0.6\}$ for CAMELS-SAM and CAMELS using the validation set. Note that $\lambda=1.0$ recovers the original LLS baseline.
Discussion of Results
As shown in Tab.˜11, all baselines (including their modifications, if applicable) perform similarly in Quijote, with notable worse performance on predicting $v_{z}$ affected by the redshift line of sight than predicting $v_{x},v_{y}$. Applying modifications in linear theory and LLS boosts their performance in CAMELS-SAM and CAMELS  [^7] The best shrinkage factor $\lambda$ for LLS modification is found to be $0.4$ for both CAMELS-SAM and CAMELS.. Surprisingly, GNN performs better in CAMELS-SAM and CAMELS when replacing real-space positions with redshift positions, due to the improvement along the line-of-sight $z$-axis. This arises from a simplified learning objective. Recall the redshift position input is a “noisy” version of the output, given by $\tilde{z}=z\cdot(1+\frac{v_{z}}{100})$. Therefore the GNN can learn to solve a simpler denoising task, rather than fitting an arbitrary map from positions (input) to velocities (output). This is supported by our ablation experiment, where we choose the line-of-sight direction as the $x$-axis and train the same GNN model from scratch. We find per-axis $R^{2}$ in CAMELS as $R^{2}(v_{x})=0.5011,R^{2}(v_{y})=0.2155,R^{2}(v_{z})=0.2215$, demonstrating the notable boost along line-of-sight $x$-axis compared to the other axes.
Table 11: Velocity Prediction from Redshift Positions. Column header “ALL” denotes overall $R^{2}$ across the three axes, whereas $v_{x}$”, $v_{y}$” $v_{z}$” denote per-axis $R^{2}$. Numbers marked with a denote that additional “cosmological oracle” information was used.
$\mathbf{R^{2}}$ $\uparrow$ Quijote CAMELS-SAM CAMELS
All $v_{x}$ $v_{y}$ $v_{z}$ All $v_{x}$ $v_{y}$ $v_{z}$ All $v_{x}$ $v_{y}$ $v_{z}$
Linear theory oracle 0.3269 0.3629 0.3629 0.2527 0.1177 0.2401 0.2390 -0.1268 0.0615 0.1653 0.1564 -0.1467
   w/ modification 0.3269 0.3628 0.3628 0.2527 0.1621 0.2902 0.2891 -0.0938 0.0732 0.1959 0.1840 -0.1686
LLS 0.3367 0.3914 0.3912 0.2274 0.1172 0.2022 0.2029 -0.0530 0.0801 0.1265 0.1134 0.0029
   w/ modification 0.3362 0.3911 0.3910 0.2266 0.2116 0.2518 0.2523 0.1310 0.2185 0.2046 0.1891 0.2604
GNN 0.3500 0.3916 0.3924 0.2643 0.3177 0.3002 0.3042 0.3477 0.3197 0.2280 0.2176 0.5067
@@ C.1 Predicting Cosmological Parameters from Merger Trees
In this section we provide further details about the 1-Nearest Neighbor predictor introduced in Sec.˜5.1. We consider for each tree the empirical distribution across nodes of three features: the concentration $c$, the maximum velocity $v_{\text{max}}$, and the scale factor $a$ (a measure of time). Consider two empirical distributions, $\widehat{P}_{A}$ and $\widehat{P}_{B}$ of a univariate random variable $X$. The Kolmogorov-Smirnov statistic [e.g. 69] is
@@
As shown in Tab.˜5 (left), $c$ and $v_{\rm max}$ are the most informative features when inferring $\Omega_{\rm m}$, while $a$ is most informative for $\sigma_{8}$. Combining all nodes features achieves far superior accuracy. To examine the model performance further, we present scatter plots of the target cosmological parameters ($x$-axis) and the predicted parameters ($y$-axis) from (a) DeepSet and (b) GNN: Fig.˜C.1 shows the results from using the node feature scale factor $a$ only, whereas Fig.˜C.2 shows the results using all features (mass $M$, concentration $c$, halo maximum circular velocity $v_{\max}$, and scale factor $a$). We see mild improvements of GNN over DeepSet; both models are highly predictive of $\Omega_{\rm m}$ when using all features. In this work, we make use of the entire merger tree from CS-Trees to predict cosmological parameters; using a subtree for such prediction task is an interesting direction for future research (e.g. the main branch used for generative modelling tasks in [88]).
@@ C.2 Reconstructing Fine-Scale Merger Trees
[…]
@@
As reported in Tab.˜5 (right), the ML approaches perform equivalently or better than the cosmology baseline (EPS). Moreover, they require considerably less time. Nevertheless, all baselines considered still have room for improvement. As a first step, we only consider predicting the unresolved merger node labels (i.e. whether there exists a merger node or not). Another important future step is to also predict the unresolved merger node features, which would enable enhanced inference quality from incomplete or coarse-grained merger trees.